The owner of a store wants to make a 30-pound mixture of two candies to sell for $3 per pound. If one candy sells for $2.95 per pound and the other for $3.10 per pound, how many pounds of each should be used?
Answer by palanisamy(496)

Let x pounds of candy A be mixed with y pounds of candy B.
Then, total x+y = 30 ...(1)
Next, candy A sells for $2.95 per pound and the other B for $3.10 per pound,
He wants to sell 30-pound mixture of two candies for $3 per pound.
So, 2.95x+3.10y = 30*3
2.95x+3.10y = 90 ...(2)
(1)*2.95 => 2.95x+2.95y = 30*2.95 = 88.5 ...(3)
(2)-(3) => 3.10y-2.95y = 90-88.5
0.15y = 1.5
y = 1.5/0.15
y = 10
(1)=> x+10=30
x = 30-10 = 20